The core diameter of multimode step index fibre is 60 micrometer. The difference in refractive index is 0.013. The core refractive index is 1.46. Predict the number of guided modes when the operating wavelength is 0.75 micrometer

 

 

Grade:12th pass

1 Answers

Vikas Amritiya
2 years ago
n12_n22=(n1+n2)=(n1+n2/2)(n1_n2/n1)2n1

n1+n2/2n1

and

NA=n122

therfore

na=n12

v-no. or normalized frequence

v=n12_n22

No. of fibre traveling

for step index fibre

N=v2/2

for graded index fibre

N=v2/4